express as a mixed radical
a) square root 98
b) -5 square root 32
how do i solve these types of questions
thanks

Answer by checkley75(3666)

sqrt98=sqrt(2*49)=[7sqrt2]
-5sqrt32=-5sqrt(2*16)=-5*4sqrt2=[-20sqrt2]